OpenStudy (anonymous):

Does anyone know what the greatest common factor of 462, 966, 336 & 378 would be?

OpenStudy (lalaly):

42

OpenStudy (anonymous):

here is how you do it, factor each number

OpenStudy (akshay_budhkar):

yay 42!!

OpenStudy (anonymous):

you get \[462=2\times 3\times 7\times 11\] \[966=2\times 3\times 7\times 23\] \[336=2^3\times 3\times 7\] and \[378=2\times 3^2\times 7\] what they all have in common is \[2\times 3\times 7=42\]

OpenStudy (anonymous):

wow! you can just cheat! http://www.wolframalpha.com/input/?i=gcf+%28+462%2C+966%2C+336%2C+378+%29

OpenStudy (anonymous):

even factors them for you. don't have to do anything at all. yikes

OpenStudy (akshay_budhkar):

legand dont teach them that..lol

OpenStudy (anonymous):

yeah well how did you do it so fast? i can't factor like that in my head, can you?

OpenStudy (akshay_budhkar):

no i cant make it very fast but by vedic maths i make at an ok rate.. i also give answers from wolfram many tims.. lol